OpenStudy (anonymous):

Hi! I need to know how to add things like (pi/3)+(pi/2). I have no idea what that type of operation is called, but I need to it to graph sin and cos.

jimthompson5910 (jim_thompson5910):

Ignore the pi's Think of it as 1/3 + 1/2

jimthompson5910 (jim_thompson5910):

How would you add those two fractions?

OpenStudy (anonymous):

find a common denominator, add the numerators, then simplify if i can

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

So it would basically be 5/6?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

1/2 + 1/3 = 5/6

OpenStudy (anonymous):

How would you put the pi back in?

jimthompson5910 (jim_thompson5910):

pi/2 + pi/3 = 5pi/6

OpenStudy (anonymous):

Thank you so much! You just made my life soooooo much easier, both of you

jimthompson5910 (jim_thompson5910):

Here's a more detailed break down pi/2 + pi/3 (3/3)*(pi/2) + (2/2)*(pi/3) 3pi/6 + 2pi/6 (3pi+2pi)/6 5pi/6 But I'm sure you see it by now

OpenStudy (anonymous):

Yep, I do, but the breakdown helped any doubts I had. Thank you for taking the time to do that, many people wouldn't, but I really appreciate it.

jimthompson5910 (jim_thompson5910):

you're welcome